OpenStudy (anonymous):

A flag pole is 12 feet tall. It's shadow is 10 feet long. What is the approximate distance from the top of the flag pole to the tip of the pole's shadow? A. 22 feet B. 16 feet C. 11 feet D. 7 feet

OpenStudy (anonymous):

XDD

OpenStudy (anonymous):

@wolf1728 can you help me please?

OpenStudy (dumbcow):

hypotenuse of a right triangle a^2 + b^2 = c^2 sound familiar?

OpenStudy (anonymous):

It's pathagreom therom i just forgot how to do it

OpenStudy (dumbcow):

yep a = 10 b=12 solve for c

OpenStudy (anonymous):

A?

OpenStudy (dumbcow):

dont guess, try to figure it out first then look at the answers plug in 10 for a , 12 for b 10^2 + 12^2 = c^2

OpenStudy (anonymous):

i got 12

OpenStudy (dumbcow):

ok what did you get for 10^2 ?

OpenStudy (anonymous):

100

OpenStudy (dumbcow):

12^2

OpenStudy (anonymous):

144

OpenStudy (dumbcow):

good, then 100+ 144 is what?

OpenStudy (anonymous):

244

OpenStudy (dumbcow):

c^2 = 244 then what did you do? not sure how you came up with 12

OpenStudy (anonymous):

nm it's B

OpenStudy (dumbcow):

good :)